OpenStudy (anonymous):

Explain how to find the equation for a vertical asymptote when given the equation.

OpenStudy (anonymous):

rational functions

OpenStudy (callisto):

Put the denominator =0 Vertical asymptotes are at the values which make the denominator zero, that is x = ..

OpenStudy (anonymous):

ok, what about discontinuities?

OpenStudy (callisto):

Hmm.. what do you mean?

OpenStudy (anonymous):

explain how you find the location of a removable discontinuity(hole) in a graph when given the equation.

OpenStudy (callisto):

I...... also put denominator =0 ...

OpenStudy (anonymous):

? both arent same though

OpenStudy (callisto):

Wait...... \[y= \frac{(x-1)(x+1)}{x-1}\]\(y= (x+1)\) where x ≠ 1

OpenStudy (anonymous):

which is that?

OpenStudy (callisto):

Then, you draw a hole at x=1 for this example. => discontinuities

OpenStudy (anonymous):

so let me get this straight, (x+1) = asymptote Whatever number that makes denominator=0 is the discontinuity?

OpenStudy (callisto):

x+1 is not the asymptote...

OpenStudy (anonymous):

other way round?

OpenStudy (callisto):

Hmm.. nope... For the function\[y= \frac{(x-1)(x+1)}{x-1}\]It can be reduced to y= (x+1). This is the one you graph. The point is the you cancel the common factor x-1 in both numerator and denominator. You need to know that you can't cancel 0 as a common factor. So, put x-1 =0, you get x=1 So, that's where the function is undefined. and the graph does not pass though the point at x=1. In other words, the graph is broken at x=1. => discontinuity

OpenStudy (anonymous):

ok, so x=1 is discont (the number that makes the denominator 0) and the vertical asymptote is?

OpenStudy (anonymous):

I understand the discontinuity part

OpenStudy (callisto):

No vertical asymptote for that example.

OpenStudy (anonymous):

well, how do i find the vertical asymptote in general?

OpenStudy (accessdenied):

A removable discontinuity occurs at the value of x that causes the function to 'equal' 0/0 A vertical asymptote occurs at the value of x that makes the denominator 0 but the numerator is still nonzero. So basically, if the numerator and the denominator share a factor of x, then there is a 'removable discontinuity' at the value of x that causes that factor to be 0, i.e. you can 'remove' it like just 'a/a' (except you still have the restriction caused by that factor in the denominator after cancelling it). If the denominator has a specific factor of x that causes only the denominator to be 0 (not in the numerator), then there is a vertical asymptote at that value.

OpenStudy (anonymous):

ok, so when i cancel, what's left is the vertical asymptote?

OpenStudy (callisto):

I fail....

OpenStudy (anonymous):

its okay, u still helped

OpenStudy (callisto):

I'm thinking... When you can cancel the common factor, put the factor you cancel =0 to get the value for discontinuity. When you see a fraction that cannot be simplified. Put the denominator =0 to get the vertical asymptote(s).

OpenStudy (accessdenied):

Yeah, that sounds good to me.

OpenStudy (callisto):

like this: \[y=\frac{x^2 -x+1}{x-1}\] x=1 is the vertical asymptote

OpenStudy (anonymous):

\[f(x)=\frac{\sum_k a_k x^k}{\prod_{j}(x-r_j)}\] implies that there is a vertical asymptote at x=r_j for all j if and only if there aren't common factors in top and bottom, which there is a "hole" or "removable" discontinuity.

OpenStudy (anonymous):

im only doing rational algebraic expressions

OpenStudy (anonymous):

Tis what that is :P

OpenStudy (anonymous):

\[\frac{a_k x^k + a_{k-1}x^{k-1}+...}{(x-r_j)(x-r_{j-1})...}\]

OpenStudy (anonymous):

?????????????????

OpenStudy (anonymous):

Alright i think i can summarize... 1. the number that makes the denominator=0 is the discont. 2. the vertical asymptote is what's left after canceling common factor plz tell me if im wrong

OpenStudy (callisto):

Ehh... the common factor cancelled =0 fir discon.

OpenStudy (anonymous):